### v1

# HLStrans: Dataset for LLM-Driven C-to-HLS Hardware Code Synthesis

Abstract
High-level synthesis (HLS) enables software developers to describe and implement hardware at a higher level of abstraction by using C/C++ instead of traditional hardware description languages to automatically generate FPGA-ready designs. However, generating HLS code significantly differs from standard C/C++: it disallows certain coding idioms, relies on specialized libraries, and critically requires fine-grained transformations and the insertion of optimization directives (pragmas) to achieve high performance. Large language models (LLMs) have shown promise in automating such transformations, yet existing open-source datasets lack sufficient complexity and optimization diversity. To address this gap, we introduce the HLStrans dataset, a comprehensive collection of 137 distinct real word programs, each annotated with a variety of C-to-HLS transformations that yield over 23K labeled design variants. These include a broad spectrum of pragmas and code-level optimizations. We benchmark state-of-the-art LLMs on this dataset to evaluate their ability to generate synthesizable, high-performance HLS code. As part of an ongoing effort, we plan to expand the HLStrans dataset in both scale and program variety, further empowering research at the intersection of AI and hardware synthesis.
1 Introduction
Specialized computing systems, also known as domain-specific architectures tailored to particular application domains, have seen dramatic growth as workloads like machine learning, signal processing, and data analytics outstrip the capabilities of general-purpose processors. By offloading compute-intensive kernels to accelerators such as Graphics Processing Units (GPUs), Field-Programmable Gate Arrays (FPGAs), or Application-Specific Integrated Circuits (ASICs), designers achieve orders-of-magnitude improvements in throughput, latency, and energy efficiency. To capitalize on this trend, begin by profiling your target workload to identify its most critical functions, then select the most appropriate hardware substrate and toolchain to implement those kernels with maximum parallelism and minimal overhead.
High-Level Synthesis (HLS) is a methodology that elevates hardware design to higher levels of abstraction by allowing engineers to describe functionality in software languages like C or C++, from which synthesizable hardware descriptions are automatically generated. The HLS tool analyzes the code, infers parallel execution paths, inserts pipeline registers, and maps data accesses to on-chip memories, producing Register-Transfer Level (RTL) descriptions in Hardware Description Languages (HDLs) such as Verilog or VHDL. Using HLS can accelerate the hardware design process and enable software engineers to engage in hardware development.
However, HLS departs radically from conventional software development by requiring developers to adopt a narrow, hardware-centric subset of C++ and acquire domain-specific expertise. Typical HLS workflows forbid features like dynamic memory allocation, unrestricted pointer arithmetic, recursion, and most of the standard library, instead demanding code that maps directly to efficient RTL ye2024hida sohrabizadeh2022autodse . To achieve high-quality output, engineers must refactor their designs so that loops have compile-time bounds, arrays possess fixed sizes, and control logic forms strictly nested or single-producer/single-consumer structures sudvarg2024hls martin2009high ye2022scalehls . On top of this, HLS tools expose a suite of vendor-specific pragma directives for loop pipelining, unrolling, array partitioning, and dataflow streaming that developers weave into their code. Many also integrate specialized C++ libraries offering arbitrary-precision fixed-point types, hardware-friendly math routines, FIFO streams, and task-oriented dataflow constructs to bridge the gap between software abstractions and synthesizable hardware modules.
LLMs offer hope to aid designers in writing HLS code. Rather than just producing HLS code from natural language with LLMs, more recent works focus on converting existing C kernels into optimized HLS implementations collini2024c2hlsc xiong2024hlspilot bhattacharyya2024llm xu2024automated prakriya2025lift , leveraging structured inputs to yield high-performance designs rather than naive, resource-heavy translations. Despite these advances, integrating LLMs into HLS toolchains remains challenging and relatively unexplored, in large part because there is no comprehensive, open-source corpus of diverse, high-quality HLS kernels. This gap prevents practitioners and researchers from fully exploiting the code-generation and optimization capabilities of modern LLMs within their design workflows.
To address the scarcity of high-quality open-source HLS benchmarks, we introduce the HLStrans dataset, a comprehensive suite of 23K HLS kernels harvested from academic repositories. Each entry pairs a naive, functionally equivalent C implementation with multiple performance-tuned HLS variants. To assemble this corpus, we developed an automated augmentation pipeline that leverages Monte Carlo tree search (MCTS) guided by electronic design automation (EDA) tool feedback and design space exploration (DSE) tools, ensuring diverse yet compiler-friendly code transformations. Every kernel is annotated with cycle-accurate latency and detailed resource-utilization metrics obtained via synthesis, creating a rich training signal for both functional correctness and performance optimization. Building on this resource, we further fine-tune large language models using a novel reinforcement-style scheme of "performance tags" and "resource tags" (rated 1–10) that explicitly reward efficient hardware design choices.
In summary, we have the following contributions:
  • HLStrans Dataset: We introduce HLStrans, the first large-scale dataset of over 23K paired C and HLS programs for LLM-powered HLS code synthesis from software.
  • LLM-Driven HLS Data Augmentation: We propose the first effective augmentation framework that leverages LLMs, MCTS, EDA feedback, and DSE tools to generate high-quality synthetic HLS code.
  • Benchmark-Driven LLM Evaluation: Enabled by our benchmark suite, we evaluate various prompts and fine-tuning methods and demonstrate that while off-the-shelf code LLMs underperform on HLS tasks, retrieval augmentation and task-specific fine-tuning on HLStrans yield substantial improvements.
2 Background
Generating high-performance HLS code from standard C for FPGAs requires a series of systematic transformations to expose parallelism, optimize data movement, and conform to HLS-friendly coding styles. While the detailed transformations are in Appendix A.1, these transformations fall into five broad categories overall:
(a) Code restructuring: loop tiling examples
(b) Directive insertion: unroll and array partition
(c) Data-type adaptation: float to fixed_type
(d) Function transformation: sqrt example
Figure 1: HLS code optimization examples
(1) HLS Code Optimization: Code Restructuring. Refactor algorithms to expose pipelining and dataflow, apply loop tiling, memory coalescing, ping-pong buffering, and reorganize control logic for parallel or streaming execution. Figure 1(a) shows an example of loop tiling.
(2) HLS Code Optimization: Directive (Pragma) Insertion. Place HLS pragmas to guide the tool scheduler, such as data flow, pipeline, loop partition, and interface specifications, to fine-tune performance and resource usage. Figure 1(b) shows an example of pragma use for unrolling and array partitioning.
(3) HLS Code Optimization: Data-Type Adaptation. Replace generic C types with HLS-specific arbitrary-precision types: convert floating point to fixed point (ap_fixed) for resource optimization, standard integers to bit-accurate (ap_int/ap_uint), and customize bit widths to match application precision requirements. Figure 1(c) shows an example of transforming float to fixed types.
(4) HLS Code Optimization: Transformation of Functions. Transform standard C functions into HLS-optimized kernels or intrinsics (such as converting the std::sqrt function to the hls::sqrt function) to better leverage FPGA fabric and specialized accelerators. Figure 1(d) shows an example of transforming the standard sqrt function to the sqrt function of the HLS library.
(5) HLS Code Repair: HLS-Compliant Coding Style. Eliminate unsupported C constructs such as dynamic memory allocation (malloc/free), recursion, and certain pointer arithmetic patterns; restructure code to use static arrays, simple loops, and explicit handshaking for communication. Figure 2 shows an example of removing recursion, which cannot be synthesized successfully.
Figure 2: HLS Code repair example: recursion elimination
3 Related works
There is an increasing body of literature on applying LLMs to generate HLS design from original C code. C2HLS collini2024c2hlsc evaluates the basic task of translating naive C++ into synthesizable HLS C++. bhattacharyya2024llm demonstrates that LLMs can automate HLS pragmas and optimizations to produce synthesizable, high-performance RTL from C on image-processing benchmarks. LIFT prakriya2025lift introduces LLM–GNN hybrid methods that automatically optimize C/C++ designs to boost FPGA performance. xu2024automated presents an LLM-driven HLS program-repair framework that transforms C/C++ into synthesizable HLS-C. HLSPilot xiong2024hlspilot extends this approach with retrieval-augmented generation and chain-of-thought prompting to deliver optimized HLS implementations across nine applications. However, to date, no work has established a large-scale dataset for thoroughly evaluating an LLM’s HLS capabilities.
First, prior HLS datasets have primarily focused on resource estimation rather than on generating HLS code from C, and available program sources remain scarce. HLSsyn bai2023towards concentrates on encompassing a diverse range of optimization pragmas into only 42 kernels for training and evaluation of design quality prediction models. HLSDataset wei2023hlsdataset , including 34 data sources and 18K samples, targets power estimates, resource, and timing estimation. MLSBench goswami2022mlsbench is an open-source dataset created with the Xilinx Vivado HLS flow, covering 17 C/C++ and 13 SystemC benchmarks drawn from Polybench, CHStone, and MachSuite. It includes only HLS log files and reports. Spector gautier2016spector comprises a collection of OpenCL applications that the authors process with the Intel OpenCL SDK to produce 8,300 distinct hardware implementations for Intel FPGAs. Dai et al. dai2018fast have open-sourced a dataset of about 1,300 designs generated by using applications from suites such as CHStone, Machsuite, and Rosetta. Second, existing datasets inadequately capture comprehensive C-to-HLS transformations, focusing largely on pragma insertion. Third, they are not organized as paired C-and-HLS examples and omit the corresponding testbenches needed for LLM-based HLS code optimization, which are not ready for LLM.
Compared with previous works, Table 1 concludes that our dataset has more kinds of sources and supports more transformations, making it ready for LLM code generation.
Table 1: Comparison of different HLS dataset
Work Samples Sources Platform & Abstraction level Supported transformation Use Case in ML Ready for LLM
Daidai2018fast 1,300 65 FPGA HLS (2) Quality of Results Estimation No
MLSBenchgoswami2022mlsbench 6,000 30 FPGA HLS (2) No
Spectorgautier2016spector 8,300 9 FPGA HLS (2) No
HLSdatasetwei2023hlsdataset 18,876 34 FPGA HLS (1) (2) (3) (4) Quality of Results Estimation. No
HLSsynbai2023towards 42,000 42 FPGA HLS (2) Quality of Results Estimation. No
Ours 23,000 137 FPGA HLS (1)(2)(3)(4)(5) Code generation. Yes
4 High-Level Synthesis (HLS) Datasets Construction
Open-source HLS datasets are scarce and unstructured, making them insufficient for training LLMs. To address this, we will develop an automated pipeline that generates high-quality HLS datasets from the limited available resources. The datasets are released under the MG0-2.0 license (duan2024modelgo, ). [^1] https://www.modelgo.li/. To construct a high-quality dataset, we need two steps. The first step is to collect high-quality open-source HLS examples. However, human-optimized code examples are not enough. The second step is the effective data augmentation for human-optimized codes to produce effective codes. Figure 3 shows our dataset construction process.
4.1 Dataset Collection
Manually optimized codebases often suffer from inconsistent formatting and minimal documentation, making them ill-suited for LLM-driven code generation. Public kernels frequently rely on unexpanded macros and include extraneous utility functions that obscure core algorithms. Our goal is to prepare C-to-HLS designs for seamless LLM consumption by packaging each design with the following files:
  • Single original file $x$ that is a slow original C/C++ codes.
  • Single optimized HLS file $y$ that implements the kernel, including a top function and, if necessary, any sub-functions and specialized data types. The file must be synthesizable and not exceed the resources of the platform.
  • Self-contained C++ testbench $tb$ includes all test cases and validation logic necessary to verify the kernel’s outputs against expected results. We manually write the testbenches and adjust the optimized HLS code to ensure it passes all tests.
Therefore, we construct pairs $(x,y,tb)$. If the original file $x$ is synthesisable, the execution cycles from synthesis reports of $y$ must be less than $y$. If the original file $x$ is not synthesisable, $y$ should be synthesisable. Table 2 shows our HLS dataset collection including common heterogeneous benchmarks such as polybench, and other FPGA benchmarks.
Table 2: Dataset collection summary
Data Source Applications Average Line of codes Supported transformations
Polybenchpouchet2012polyhedral 28 23 (1)(2)(3)
Rodiniacong2018understanding 15 61 (1)(2)(3)
Machsuitereagen2014machsuite 16 59 (2)(3)
Rosettazhou2018rosetta 3 127 (1)(2)(3)(4)
C2HLScollini2024c2hlsc 14 47 (2)(4)(5)
PP4FPGAkastner2018parallel 4 36 (1)(2)(3)(4)
Forgebenchwanna2025forgebench 6 112 (4)
HLSfactoryabi2024hlsfactory 11 107 (1)(2)(3)(4)
Others (GitHub) 41 101 (1)(2)(3)(4)(5)
Totals 137 93 (1)(2)(3)(4)(5)
4.2 Dataset Augmentation
Given the performance, resource utilization, and compilation feedback provided by EDA toolchains, we employ Monte Carlo tree search (MCTS) and design space exploration (DSE) tools to efficiently explore the large and combinatorial action space as shown in Figure 3.
First, MCTS performs structured exploration by balancing the exploitation of high-reward actions with the exploration of uncertain or less-visited regions of the search space. The optimization policy is generated by the retrieval augmentation generation (RAG) module. The search is guided to choose the suitable policy by both the verification pipeline and a reward model. The reward model incorporates detailed feedback from the HLS toolchain, including synthesis success or failure, compile warnings, and performance metrics such as resource usage, latency, and throughput. This heuristic-driven strategy enables the agent to iteratively refine transformation sequences and produce more high-quality, synthesizable HLS designs.
Second, HLS directive design space exploration using genetic algorithms ferikoglou2023collectivehls is adopted. It inserts pipeline, unroll, and partition pragmas to produce more effective data samples. Through iterative refinement, the framework converges toward optimized and synthesizable HLS code. After collecting these C-to-HLS code pairs, we use an LLM to produce a chain-of-thought narrative that analyzes the performance and resource bottlenecks, explains why the original code cannot be synthesized, and highlights the key optimizations applied in the HLS version. This explanatory reasoning is then integrated alongside the optimized HLS output.
Figure 3: HLStrans Dataset Augmentation Framework
4.2.1 Monte Carlo tree search (MCTS)
MCTS enables an agent to learn to navigate the vast space of possible code transformations while balancing multiple optimization objectives. The agent’s decisions are guided by comprehensive feedback from the HLS toolchain, including synthesis success, resource utilization, and performance metrics.Our MCTS has the following elements:
Environment E: The environment is composed of the HLS toolchain, specifically Xilinx Vitis HLS, which compiles the code and provides critical feedback such as performance estimates.
Agent G: We propose to use an LLM as the agent that leverages its pretrained knowledge of hardware design and in-context learning abilities.
Action A: At each time step $t$, the agent selects an action $a_{t}$, which corresponds to a prompt or transformation applied to the current HLS code. We define two complementary action types: RAG-based actions retrieve optimization policies directly from our pre-built table and accompanying code examples shown in Figure 4, leveraging retrieval-augmented generation to surface proven transformations rapidly and reliably. Reasoning-based actions with ReAct prompt yao2023react , in contrast, analyze compiler warnings such as pipeline-interval breaches or loop-unroll violations and apply targeted code reforms by interpreting warning semantics within the current code context.
(a) Loop tiling Code examples by RAG
(b) LLM reasoning about environment warning and tool hint
Figure 4: Actions design of MCTS
State S: The state $S_{t}$ at time step $t$ is defined as the current version of the HLS code after applying the previous actions.
Reward R: Rule-based reward shaping has proven effective in guiding agent behavior in previous work guo2025deepseek . In our framework, the reward function $R(s_{t},a_{t})$ is computed by applying rule-based scoring to verification results and feedback provided by the HLS tool. A penalty of $-2$ is applied when the verification fails, and $-1$ if synthesis fails or the design exceeds resource constraints. A neutral reward of 00 is given when the transformed design performs worse than the original, while a reward of $1$ is granted when it performs better. If the design not only surpasses the original but also meets timing constraints, a higher reward of $2$ is assigned.
In our cases, the MCTS begins at the initial state $S_{0}$ (the root node), which is the naive HLS code. From a state $S_{t}$, the agent applies an optimization policy $\pi$, i.e., an action $a_{t}\in\mathcal{A}$, transitioning to the subsequent state $S_{t+1}$. MCTS consists of four key phases: selection, expansion, simulation, and backpropagation.
Selection: We employ the upper confidence bounds for trees (UCT) gelly2006exploration algorithm to choose nodes.
$\pi(s_{t})=\arg\max_{a_{t}\in\mathcal{A}}\left(\underbrace{R(s_{t},a_{t})}_{\text{reward}}+\beta\times\underbrace{\frac{\sqrt{1+N(s_{t})}}{1+N(s_{t},a_{t})}}_{\text{ $U$ Term}}\right).$
Expansion: From the current state $s_{t}$, generate one or more child nodes to explore untried actions, formalized as $a_{t+1}=\mathrm{llm}(s_{t})$.
Simulation: Perform a rollout from the chosen child node by applying $a_{t+1}$, running HLS synthesis to estimate latency and resource utilization, and computing the reward $R(s_{t},a_{t+1})$.
Backpropagation: Propagate the obtained reward back up the visited path, updating each node’s statistics (e.g., visit count and value estimate) to improve future selection.
4.2.2 Design Space Exploration
The tool implements an automated HLS design-space explorer that uses a genetic-algorithm optimizer to discover effective directive combinations, specifically loop pipelining, loop unrolling, and array partitioning, that maximize performance and resource utilization. By encoding each candidate solution as a set of directive parameters, evolving them through mutation and crossover, and evaluating them via HLS synthesis feedback, the tool generates high-quality configurations without manual tuning.
Overall, we leverage DeepSeek-R1 to generate high-quality synthetic code examples, the AMD Vitis HLS EDA tool, and DSE tools to validate, annotate, and collect performance/resource metrics within our framework, yielding an effective, end-to-end HLS code dataset. The detailed information on dataset generation is in Appendix A.2.
5 Experimental Evaluation of LLMs on Our Dataset
5.1 Prompting Methods
Zero-shot Prompting: We craft concise, HLS-specific prompts that instruct the model to perform code optimization or transformation from its pretrained knowledge, without any additional fine-tuning or example demonstrations liu2021makes wei2021finetuned .
Chain-of-Thought Prompting: Building on the chain-of-thought approach of Wei et al. wei2022chain , our prompts first guide the model through a transformation reasoning phase before asking it to emit the refined code.
Retrieval-Based Prompting: Recent studies shrivastava2023repository shypula2023learning have shown that retrieval-based techniques can substantially boost code generation quality in large language models. In our approach, we first encode each program using CodeBERT zhou2023codebertscore to produce rich, semantically informed embeddings. We then index these vectors with FAISS johnson2019billion (Facebook AI Similarity Search) and perform a K-nearest-neighbors lookup to retrieve the top K most similar code snippets from our training corpus. Finally, we supply these retrieved examples alongside the original code as additional context to the LLM, guiding it to produce more accurate and effective edits. In our experiments, we set K to 1. The detailed prompt information is in Appendix A.3.
5.2 Experiment setting
Fine tuning setting. Training was conducted using 2 NVIDIA H100 GPUs, each with 80 GB of memory. The computing environment was configured with CUDA 12.2 and cuDNN 9.1 to ensure optimal deep learning performance.
Optimization Conditioned Fine-tuning: Programs that solve the same task can vary widely in how well they perform because of many different kinds of directives and code transformations. The optimized code has a jumble of small or large speedups, without any signal about which gains are most significant. To give the model clearer guidance, we borrow ideas from shypula2023learning and attach a “performance tag” and "resource tag" to each solution during training. Each tag reflects how close that program comes to the best attainable performance with resources minimized, using a discrete scale from 1 to 10, respectively. Concretely, for every benchmark problem, we rank its optimized solutions by latency and resource, labeling the top 10% in performance as a “10/10” performance tag and the lowest 10% in resource usage as a “10/10” resource tag. By examining these tags, the model learns which code patterns tend to yield the highest performance. At inference time, we prompt the model to generate the best-performing codes with the least resources. The detailed fine-tuning information is in Appendix A.4.
Models. We evaluate closed-source models, GPT-4o and deepseek-R1. We also used pretrained checkpoints of open-source Qwen models yang2024qwen2 : Qwen2.5-Coder-32B-Instruct, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-3B-Instruct obtained via HuggingFace wolf2020transformers .
Metrics. To evaluate the effects of HLS code generations, we use the following metrics:
  • Functional Accuracy (%Function Correct): The share of test programs that preserve the original functionality testbench.
  • Synthesis Accuracy (%Synthesis Correct): The share of programs that can be synthesized.
  • Reduce Latency (Acceleration): The ratio of original latency to new latency, $accel=o/n$, where $o$ is the old latency and $n$ is the new latency from reports of HLS synthesis respectively. We report the minimum, average, and maximum acceleration only for the functionally and syntactically correct cases.
  • Optimization Rate (%OPT): The percentage of test programs that reduce latency (Acceleration > 1) while remaining functionally correct.
FPGA Flow Setting: We transform the given sequential C++ code of these benchmarks into optimized HLS kernel code. We adopt the Vitis HLS design flow, using the Xilinx Alveo U55C data center accelerator card. The target frequency is 200 MHz.
We selected 17 applications as test cases both unsynthesizable designs requiring repair and synthesizable designs in need of optimization and used the remaining 120 applications for training (90%) and validation (10%). Previous works li2022competition show that generating multiple program candidates per input and selecting the optimal one improves code synthesis performance. We generate $k$ program variants for each input, then select the fastest one that successfully passes all test cases; we refer to this sampling-and-selection strategy as $Best@k$.
5.3 Experiment Results
5.3.1 Results of Pretrained Models
Table 3 and 4 report Best@1 and Best@5 accuracy for different prompts. Overall, deepseek-R1 achieves the highest performance, and every model benefits from chain-of-thought prompts compared to zero-shot. However, we observe that some LLM-optimized codes actually degrade performance. This can happen for two reasons: First, the restructuring performed by the LLM can introduce new loop dependencies, increasing latency; Second, the pragmas inserted by LLMs may be less effective than the default optimizations inferred by the HLS compiler.
Table 3: Best@1 Results for various methods and models.
Acceleration
Method Model Opt Min Avg Max Function correct Synthesis correct
Zero-shot Deepseek-R1 17.6% 0.3× 2.8× 24.6× 47.1% 29.4%
GPT-4o 17.6% 0.5× 1.6× 10.9× 35.3% 29.4%
Qwen coder 32B 5.9% 0.8× 1.3× 1.5× 41.2% 11.8%
Qwen coder 7B 0% 0.66× 0.9× 1.0× 38.8% 11.8%
Qwen coder 3B 0% 0.75× 0.8× 1.0× 11.8% 11.8%
COT Deepseek-R1 17.6% 0.9× 4.3× 41.0× 52.9% 29.4%
GPT-4o 29.4% 0.5× 1.3× 4.0× 52.9% 35.3%
Qwen coder 32B 11.8% 0.1× 2.4× 24.6× 58.8% 17.6%
Qwen coder 7B 17.6% 0.04× 1.2× 2.0× 52.9% 35.3%
Qwen coder 3B 5.9% 0.8× 1.1× 1.8× 47.1% 29.4%
Retrieval Prompt Deepseek-R1 41.2% 0.9× 7.2× 57.0× 47.1% 47.1%
GPT-4o 41.2% 0.9× 4.7× 24.6× 58.8% 47.1%
Qwen coder 32B 29.4% 0.7× 4.9× 24.6× 47.1% 35.3%
Qwen coder 7B 17.6% 0.3× 2.1× 5.1× 52.9% 35.3%
Qwen coder 3B 11.7% 0.7× 1.0× 3.4× 29.4% 11.8%
Retrieval-augmented prompting consistently outperforms purely generative methods, achieving the highest synthesis success rates without degrading functional correctness. This demonstrates that our dataset by providing a rich cache of validated pragmas and code transformation examples serves as an indispensable “best practices” repository, steering LLMs toward hardware-friendly idioms and dramatically reducing synthesis failures.
Table 4: Best@5 Results for various methods and models.
Acceleration
Method Model Opt Min Avg Max Function correct Synthesis correct
Zero-shot Deepseek-R1 23.5% 0.8× 4.3× 23.5× 47.1% 29.4%
GPT-4o 23.5% 0.7× 2.4× 23.5× 41.1% 29.4%
Qwen coder 32B 6.9% 0.9× 2.1× 20.4× 64.7% 35.3%
Qwen coder 7B 0% 0.05× 0.5× 1.0× 47.1% 17.6%
Qwen coder 3B 5.9% 0.59× 1.4× 8.0× 52.9% 23.5%
COT Deepseek-R1 23.5% 0.5× 8.2× 53.2× 52.9% 41.2%
GPT-4o 29.4% 0.7× 3.1× 17.1× 52.9% 35.3%
Qwen coder 32B 17.6% 0.5× 3.5× 20.4× 52.9% 35.3%
Qwen coder 7B 17.6% 0.1× 0.7× 1.0× 47.1% 23.5%
Qwen coder 3B 11.8% 0.6× 1.5× 2.3× 47.1% 35.3%
Retrieval Prompt Deepseek-R1 47.2% 0.9× 9.4× 83.2× 64.7% 58.8%
GPT-4o 41.2% 0.6× 5.9× 37.3× 58.8% 47.1%
Qwen coder 32B 35.3% 1.2× 5.1× 24.6× 47.1% 35.3%
Qwen coder 7B 11.8% 0.5× 2.7× 5.1× 52.9% 11.8%
Qwen coder 3B 17.6% 0.6× 1.5× 3.2× 35.3% 29.4%
5.3.2 Results of Fine-Tuning Models
Both the 3B and 7B fine-tuned models show consistent gains in optimization quality, latency reduction, and synthesis success rate in Table 5. They generate HLS code that not only executes faster but also synthesizes more reliably, even though the function-correct rate has slightly dropped. These results demonstrate that training on our curated dataset significantly boosts an LLM’s ability to produce correct, high-performance HLS implementations directly from C sources.
Table 5: Fine-Tuning: Best@1 vs. Best@5 Results for Various Models
Acceleration
Setting Model Opt Min Avg Max Function correct Synthesis correct
Best@1 Qwen coder 3B 11.8% 0.1× 2.1× 7.7× 35.3% 35.3%
Best@1 Qwen coder 7B 17.6% 0.4× 2.8× 17.2× 41.1% 29.4%
Best@5 Qwen coder 3B 23.5% 0.6× 2.9× 8.9× 41.1% 35.3%
Best@5 Qwen coder 7B 29.4% 0.8× 3.1× 21.4× 41.1% 35.3%
6 Conclusion
We introduce a novel dataset that transforms C kernels into richly annotated HLS implementations, empowering LLMs to learn hardware-aware optimizations such as loop pipelining, unrolling, and memory buffering. Our experiments demonstrate that retrieval and fine-tuning on this dataset significantly boosts both latency reduction and synthesis success rates, proving its effectiveness in accelerating and automating electronic design flows. By releasing both the dataset and training scripts, we aim to catalyze further exploration at the intersection of LLMs and hardware design.
7 Future work
We evaluate HLS designs based on synthesis reports, which yield resource and timing estimates rather than true hardware performance. In future work, we will measure and compare real execution time across different models on physical FPGA platforms.
Unlike large software-optimization datasets, publicly available HLS benchmarks remain scarce. To address this gap, we will explore more effective methods for generating HLS-specific datasets and continue to grow our collection of real-world examples. Our current dataset targets AMD FPGAs; we plan to extend this work by generating analogous datasets for Intel FPGAs and Catapult HLS.
Appendix A Appendices
A.1 HLS Code Transformations
A.1.1 HLS Code Optimization: Code Restructuring.
In our datasets, we apply a suite of code-reconstruction techniques designed to optimize memory access patterns, alleviate computational bottlenecks, and resolve loop dependencies. By restructuring data flow and exploiting hardware parallelism, these methods boost throughput and shorten overall execution time. Table 6 list the main Code Restructuring adopted in our dataset.
Table 6: HLS Code Reconstruction Methods
Optimization Explanation Performance Benefit
Memory coalescing Merge multiple memory accesses into one memory transaction. Reduces memory access latency and improves bandwidth utilization.
Local tiling Divide loops into tiles to improve cache reuse and spatial locality. Enhances data locality and on-chip buffer efficiency.
Ping pong buffer Alternate between two buffers for simultaneous load and compute. Hides memory latency by overlapping computation with memory access.
Dataflow Separate tasks into pipeline stages for concurrent execution. Allows function-level parallelism, boosting throughput.
Control flow optimization Replace if–else with ternary or simplified logic conditions. Reduces combinational path length, improving timing and synthesis.
A.1.2 HLS Code Optimization: HLS Directive (Pragma) Insertion.
Our dataset features an extensive catalog of HLS pragmas ranging from memory-access directives (array partitioning, streaming) through loop-level transformations (unrolling, merging, tiling) to fine-grained pipeline controls (initiation interval tuning, dataflow regions). By systematically applying and combining these pragmas, these directives empower automated HLS flows to tailor synthesized hardware for domain-specific latency, throughput, and area requirements making our dataset a valuable reference for exploring pragma-driven performance tuning. Table 7 introduce these applied pragma optimizations.
Table 7: HLS Directive (Pragma) Insertion Methods
Optimization Explanation Pragma Example
Array partition Split a large array into multiple smaller memories #pragma HLS ARRAY_PARTITION variable=arr complete
Memory type Specify the on-chip storage type (BRAM/URAM/SMALL_RAM) #pragma HLS RESOURCE variable=buf core=RAM_2P
Loop unroll Replicate loop body to create parallel compute units #pragma HLS UNROLL factor=4
Loop merge Merge consecutive loops to reduce control overhead #pragma HLS LOOP_MERGE
Function inline Inline functions to eliminate call overhead #pragma HLS INLINE
Pipeline Pipeline loops or functions to lower initiation interval #pragma HLS PIPELINE II=1
Dataflow Enable task-level parallelism between functions #pragma HLS DATAFLOW
Dependence Declare data dependencies to allow safe loop optimizations #pragma HLS DEPENDENCE variable=arr inter false
Stream Use streaming interfaces to transfer data via FIFOs #pragma HLS STREAM variable=fifo depth=8
A.1.3 HLS Code Optimization: Data-Type Adaptation.
Our dataset also incorporates a comprehensive suite of data-type adaptations optimized for FPGA synthesis. We translate generic C types (e.g., int, float, struct) into precise HLS constructs such as ap_uint<W>, ap_fixed<TOTAL,INT>, and hls::stream<T> to fully exploit on-chip LUT/FF and DSP resources. These mappings enable fine-grained control over bit-width, data alignment, and streaming interfaces, ensuring maximal throughput, minimal logic utilization, and lower power consumption in FPGA deployments. Table 8 lists the specific datatype conversions applied in our framework.
Table 8: Adaptation of C Data Types to HLS Data Types
Original C Type HLS Type Purpose
int/short/char ap_uint<$W$>/ ap_int<$W$> Precisely control integer bit-width to save LUT/FF resources
float/double ap_fixed<TOTAL,INT>/ ap_ufixed<TOTAL,INT> Replace floating-point with fixed-point to reduce DSP usage and power
struct/union struct { ap_uint<…> field; … } with bitfields Precisely specify field bit-widths and alignment, eliminate padding
pointer/array hls::stream<T> Map to hardware FIFO streams for stream transmission
A.1.4 HLS Code Optimization: Transformation of Functions
By transforming standard C/C++ functions into their corresponding HLS intrinsics, developers can leverage highly optimized FPGA kernels. This approach dramatically boosts execution performance by exploiting dedicated hardware units for math operations and data manipulation. At the same time, it conserves FPGA resources, reducing logic utilization and power consumption compared to generic software approximations. Table 9 list the transformations of standard math functions to HLS intrinsics.
Table 9: Transformation of Standard Functions to HLS Intrinsics
Standard C/C++ Function HLS Intrinsic Purpose
std::sqrt(x) hls::sqrt(x) Generates a pipelined square root unit instead of slow software approximation
std::exp(x) hls::exp(x) Synthesizes an exponential function hardware block (LUT-based)
std::log(x) hls::log(x) Provides a hardware-friendly implementation of natural logarithm
std::sin(x) hls::sin(x) Efficient sine computation using CORDIC or LUTs
std::cos(x) hls::cos(x) Efficient cosine computation using CORDIC or LUTs
a / b hls::div(a, b) Replaces division with a synthesizable divider core
a % b hls::mod(a, b) Synthesizes modulo operation in hardware
A.1.5 HLS Code Repair: HLS-Compliant Coding Style.
High-level synthesis (HLS) cannot synthesize all idiomatic C constructs directly. To enable hardware generation, we must refactor unsupported patterns like dynamic memory allocation, recursion, and pointer arithmetic into HLS-compliant coding styles that the tool can analyze and map to on-chip resources. Table 10 lists these common transformations.
Table 10: Transformation of Unsupported C Constructs for HLS Compatibility
Unsupported C Construct Recommended HLS-Compatible Transformation Purpose
Dynamic memory allocations Use static arrays with fixed size at compile time HLS tools require compile-time memory size to synthesize physical resources (BRAM/LUTRAM)
Recursion Convert to iterative form using for/while loops Recursion creates a dynamic call stack, which is not synthesizable
Pointer arithmetic beyond array indexing Use bounded array indexing Allows compiler to infer memory access patterns and pipeline-optimize
Function pointers or callbacks Inline or manually instantiate function variants HLS requires all control flow to be static and analyzable at compile time
Variable-length arrays Replace with fixed-size arrays defined by constants or macros HLS cannot synthesize dynamically sized buffers
A.2 Dataset generations
A.2.1 Dataset collections
We began by collecting a comprehensive set of human-optimized HLS implementations from GitHub repositories and real-world FPGA applications. Our curated corpus spans diverse application domains and covers all five transformation categories detailed in Appendix A.1. Figure 5 visualizes the distribution of programs across these five tasks, and the source kernels themselves fall into seven distinct application categories. This rich, well-balanced dataset provides the broad coverage of real-world HLS patterns necessary to train and evaluate LLM’s hardware synthesis ability.
Figure 5: dataset program source distribution
A.2.2 Dataset Augmentation
Retrieval-Augmented Generation: To broaden the range of HLS code-transformation techniques that our LLM can learn, we built an automated framework (see Appendix A.2) that programmatically generates optimized variants via Monte Carlo Tree Search. Central to this system is a Retrieval-Augmented Generation (RAG) table of optimization strategies including code-reconstruction patterns, directive (pragma) insertions, data-type adaptations, and function-level transformations each entry pairing a concise description with a few-shot example illustrating the baseline code and its optimized counterpart listed in A.1. During search, these RAG-driven actions guide the MCTS policy to apply specific transformations, yielding a diverse corpus of HLS kernels ready for LLM fine-tuning and evaluation. One kind of Retrieval-Augmented strategies is shown in Fig.6 and prompt template is shown in Fig.7.
Figure 6: Example of Retrieval-Augmented strategies in MCTS framework
Figure 7: Prompt Template for the optimization with MCTS framework
Framework Evaluation: We evaluated our dataset-augmentation framework on the widely adopted Rodinia benchmark suite [38], using a Xilinx Alveo U55C FPGA board running at a 300 MHz kernel clock. Our goal was to measure how effectively our MCTS-based sampler could guide Deepseek-R1 and GPT-4o toward highly optimized HLS kernels.
Table 11 summarizes the Best@1 kernel runtimes (in milliseconds) across twelve diverse applications, comparing four configurations: Baseline (The unmodified, compiler-generated HLS implementation), HLSPilot [7] (A recent LLM based optimization framework), GPT-4o with our framework and Deepseek-R1 with our framework. Our results reveal several key findings:
  • Consistent Improvement over previous work. In every benchmark, both of our enhanced pipelines outperform HLSPilot, demonstrating that the combination of large-model code generators with MCTS exploration yields more hardware-efficient HLS designs.
  • Deepseek-R1 with our framework achieves up to average 28× reduction in real execution time compared to the baseline.GPT-4o with our framework attains up to average 20× reduction in real execution time.
  • Robust Gains Across Diverse Kernels. From compute-bound codes such as kmeans, mgvf, and streamcluster, to memory-sensitive workloads like hotspot and nw, our framework consistently identifies and applies scheduling, pipelining, and data-partitioning transformations that exploit the parallelism and memory hierarchy of the Alveo U55C.
Table 11: Runtime (ms) of different benchmarks across models.
Application Baseline HLSPilot [7] Ours
GPT-4o GPT-4o Deepseek-R1
cfd_flux 13 6.71 4.57 1.61
hotspot 1879.1 712.7 300.5 22.3
kmeans 2243.2 65.9 17.9 15.7
knn 17.0 2.8 0.83 0.82
dilate 48.8 16 0.75 1.64
gicov 107.0 93 82.3 30.7
mgvf 8047.5 3212 1231 446
lud 226.4 112 81.2 52.6
nw 206.4 145 73 13
pathfinder 7.8 5.9 1.09 1.51
srad 35.7 9.4 6.4 6.6
streamcluster 16173 9388 8162.3 3966
A.2.3 Dataset Examples
This section presents three real HLS code transformation pair examples: performance optimization (Code restructuring and Directive insertion), synthesizability correction (Code repair), and adaptation from C-style to HLS-style code (Data-type adaptation and transformation of functions). The dataset is hosted at https://huggingface.co/datasets/qingyun777yes/HLStrans.
1. Performance Optimization
Figures 8(a) and 8(b) show a simple K-Nearest Neighbors (KNN) implementation before and after HLS optimization. The optimized version achieves better performance due to improved pipelining, parallelism, and memory optimization.
(a) Unoptimized KNN implementation
(b) Optimized KNN implementation
Figure 8: Comparison of KNN implementations: (a) Unoptimized and (b) Optimized for high performance.
2. Synthesizability Transformation
Figures 9(a) and 9(b) illustrate the transformation from a non-synthesizable function into a valid HLS-compatible version.
(a) Original non-synthesizable code
(b) Modified synthesizable code
Figure 9: Transformation from non-synthesizable code to synthesizable code: (a) Original and (b) Modified version.
3. C-style to HLS-style Conversion
Figures 10(a) and 10(b) demonstrate how C-style data types and functions can be adapted into HLS-friendly forms.
(a) Original C-style code using standard data types
(b) Transformed HLS-style code using synthesizable types
Figure 10: Transformation from traditional C-style to HLS-style coding: (a) Original code and (b) Synthesizable HLS code.
A.3 Prompt Details
We explore three types of prompts used for HLS code transformation: zero-shot, chain-of-thought, and retrieval-augmented.
Figure 11: Zero-shot prompt used for HLS code transformation.
Figure 12: Chain-of-thought prompt for step-by-step transformation.
Figure 13: Retrieval-augmented prompt for enhanced transformation.
A.4 Detailed Experiment results
Unlike traditional software code, which need only pass functional correctness tests, HLS-generated kernels must also successfully synthesize and implement via the Vitis HLS toolchain to be deployable on FPGA hardware. Below, we briefly describe how we leverage synthesis results for design evaluation. Also, we introduce the fine tuning results during training.
A.4.1 HLS Synthesis Results Example
While HLS synthesis cannot yield perfectly accurate timing or resource-utilization numbers, it provides essential estimates for comparing design variants. Figures 14(a) and 14(b) show the synthesis reports for the unoptimized and optimized KNN kernels, respectively, targeting a Xilinx Alveo U55C accelerator at a 300 MHz kernel clock.
The optimized design trades increased resource usage more DSP slices, flip-flops (FFs), and lookup tables (LUTs) for a dramatic fourfold reduction in latency cycles (from 2,097,324 cycles down to 508,479 cycles). In a real FPGA deployment, this corresponds to an end-to-end runtime of approximately 3.2 ms versus roughly 17 ms for the unoptimized kernel, while still remaining within resource budgets. We report acceleration based on the estimated latency cycles from the synthesis reports. However, HLS synthesis itself can be time-consuming, particularly for large or highly optimized designs.
(a) Unoptimized KNN HLS synthesis results
(b) Optimized KNN HLS synthesis results
Figure 14: (a) Unoptimized and (b) optimized KNN implementations after HLS synthesis.
A.4.2 Finetune Results
With our real-world corpus, we reserve two C programs for the repair task and another 15 programs for the optimization task. The remaining 120 programs are split into training and validation sets. Figure 15(a) presents the validation loss curve for Qwen2.5-Coder-3B-Instruct, while Figure 15(b) shows the corresponding curve for Qwen2.5-Coder-7B-Instruct during fine-tuning. In both cases, the steadily decreasing loss demonstrates that fine-tuning effectively adapts the models to our dataset.
(a) Validation results for Qwen2.5-Coder-3B-Instruct during fine-tuning.
(b) Validation results for Qwen2.5-Coder-7B-Instruct during fine-tuning.
Figure 15: Validation performance of Qwen2.5-Coder models during fine-tuning on the held-out dataset.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims in the abstract and introduction reflect the scope of the paper and are supported by the experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes an explicit limitations in Section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not contain theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides the detailed settings to reproduce the results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will release the data and code in the non-anonymized version.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the training details, including hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Fine-tuning full datasets introduces negligible variance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report in experiment settings.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We reviewed the Code of Ethics and confirmed that this work follows the code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss in Section 7.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Details are included in the license.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer:[Yes]
  59. Justification: Details are included in the license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We discuss the documentation of data collection in detail.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not include crowdsourcing experiments and research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not include crowdsourcing experiments and research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We discuss the collection of dataset and the training methods.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.